### v5

# AsymFormer: Asymmetrical Cross-Modal Representation Learning for Mobile Platform Real-Time RGB-D Semantic Segmentation

Abstract
In the realm of robotic intelligence, achieving efficient and precise RGB-D semantic segmentation is a key cornerstone. State-of-the-art multimodal semantic segmentation methods, primarily rooted in symmetrical skeleton networks, find it challenging to harmonize computational efficiency and precision. In this work, we propose AsymFormer, a novel network for real-time RGB-D semantic segmentation, which targets the minimization of superfluous parameters by optimizing the distribution of computational resources and introduces an asymmetrical backbone to allow for the effective fusion of multimodal features. Furthermore, we explore techniques to bolster network accuracy by redefining feature selection and extracting multi-modal self-similarity features without a substantial increase in the parameter count, thereby ensuring real-time execution on robotic platforms. Additionally, a Local Attention-Guided Feature Selection (LAFS) module is used to selectively fuse features from different modalities by leveraging their dependencies. Subsequently, a Cross-Modal Attention-Guided Feature Correlation Embedding (CMA) module is introduced to further extract cross-modal representations. This method is evaluated on NYUv2 and SUNRGBD datasets, with AsymFormer demonstrating competitive results with 54.1% mIoU on NYUv2 and 49.1% mIoU on SUNRGBD. Notably, AsymFormer achieves an inference speed of 65 FPS and after implementing mixed precision quantization, it attains an impressive inference speed of 79 FPS on RTX3090. This significantly outperforms existing multi-modal methods, thereby demonstrating that AsymFormer can strike a balance between high accuracy and efficiency for RGB-D semantic segmentation. Code: https://github.com/Fourier7754/AsymFormer
I INTRODUCTION
Real-time semantic segmentation methods are important for robots as they rely on obtaining semantic information in real-time to support navigation and task decision-making. Current Real-Time semantic segmentation networks achieve competitive results and reach over 100 FPS inference speed on simple outdoor environments (Cityscapes 19 classes) [3, 4, 2, 1]. However, these networks perform poorly on complex indoor environments, such as NYU Depth v2 (40 classes) [5] and SUNRGBD (37 classes) [6]. Several studies have explored how to improve indoor scene semantic segmentation performance by integrating RGB-D information [7, 8, 9, 10], which is widely available from RGB-D sensors on indoor robots. RGB-D information consists of RGB (color, texture and shape) and Depth (boundaries and relative location) features, which are somewhat complementary [7, 8].
At present, RGB-D semantic segmentation methods have achieved state-of-the-art performance by employing dual-branch backbone for feature extraction and attention mechanisms for feature selection [8]. However, there is a dearth of discourse on whether these operations can be executed with less redundancy and higher efficiency. Primarily, most methodologies utilize two symmetrical, computationally intensive backbones, which effectively doubles the overall computational complexity [12]. Secondly, numerous methods concentrate on designing complex feature fusion modules to enhance accuracy, but do not exhibit significant advantages than some methods that employ simpler attention modules like ESANet [9], which indicates that as more computational resources are invested, the incremental improvement in accuracy decreases. Lastly, the use of cross-attention modules to jointly utilize multi-modal features has been demonstrated to be effective in improving network accuracy [8, 28]. However, most methods do not discuss how to implement cross-attention more efficiently. Therefore, the reduction of redundancy in dual-branch networks and the construction of more efficient attention modules remain significant topics for discussion.
To address this issue, this study suggests a balanced approach between computational efficiency and feature richness in multi-modal representation learning. RGB images are found to contain more information than Depth images, thus it is not necessary to use the same size backbone for extracting Depth information. The performance improvements achieved by using attention mechanisms for various tasks come with minimal increase in computational cost. While existing modules like SE [34] are efficient for feature selection, they lack the ability to model correlations in spatial dimensions. Increasing parameters can facilitate spatial dimension feature selection, but careful design is needed to maintain computational efficiency. For feature extraction, the commonly used Multi-Head Self Attention (MHSA) [22] is effective for single-modal features despite its small parameter count. Extending this efficient feature extraction to multiple modalities is feasible, but consideration should be given to feature fusion while maintaining computational efficiency.
This paper introduces AsymFormer, a high-performance real-time network for RGB-D integration semantic segmentation that employs an asymmetrical backbone design. This includes a larger parameter backbone for RGB features and a smaller backbone for the Depth branch. Regarding framework selection, Vision Transformer are always performance better, but slower than same computational complexity CNNs due to a lack of hardware optimization [40]. In order to speed up the main branch, this paper employ a hardware-friendly CNN [16] for the RGB branch and a Transformer [17] with fewer parameters but similar performance for the Depth branch to further compress the parameters. This paper also discuss feature selection modeling and construct the Local Attention-Guided Feature Selection (LASF) module. This module computes the spatial-channel attention weights in parallel to enhance inference speed. Moreover, this paper establishes a novel approach to model spatial attention. This module can estimate the differences in features from different modalities by utilizing learnable channel weights and calculating spatial attention weights for each pixel. Additionally, a Cross-Modal Attention (CMA) module is introduced to embed cross-modal integration learning information into pixel-wise fused features. Finally, we employ a lightweight MLP-Decoder[17] to decode semantic information from shallow features.
This paper evaluates AsymFormer on two classic indoor scene semantic segmentation datasets: NYUv2 and SUNRGBD. Meanwhile, the inference speed test is also preformed on Nvidia RTX 3090 platform. The AsymFormer achieves 54.1% mIoU on NYUv2 and 49.1 mIoU on SUNRGBD, with 65 FPS inference speed. After performing mixed precision quantization on the model using the TensorRT framework, AsymFormer achieves an impressive inference speed of 79 FPS on RTX 3090, which significantly surpasses existing methods. Our experiments highlight AsymFormer’s ability to acquire high accuracy and efficiency at the same time. Our main contributions are summarized as follows:
  • $\bullet$
    We employed an asymmetric backbone that compressed the parameters of the Depth feature extraction branch, thus reducing redundancy.
  • $\bullet$
    We introduced the LAFS module for feature selection, utilizing learnable feature weights to calculate spatial attention weights.
  • $\bullet$
    We introduce a novel efficient cross-modal attention (CMA) for modeling of self-similarity in multi-modal features, validating its capability to enhance network accuracy with minimal additional model parameters.
II Related Works
RGB-D Representation Learning. One of the earliest works on RGB-D semantic segmentation, FCN [35], treated RGB-D information as a single input and processed it with a single backbone. However, subsequent works have recognized the need to extract features from RGB and Depth information separately, as they have different properties. Therefore, most of them have adopted two symmetric backbones for RGB and Depth feature extraction [36, 37, 9, 11, 27, 8]. Further, the ACNet [25] proposed a three-stream backbone to process RGB, Depth and Fusion features independently. Moreover, some works have considered the variability of RGB-D features and used asymmetric backbones for their feature extraction. For example, TSNet [39] used ResNet [31] for RGB feature extraction and VGG (without residual connections) [38] for Depth feature extraction. The PSCNet [12] also used an asymmetric backbone and reduced the redundant computational cost by cutting the parameters of the Depth branch.
RGB-D feature fusion. The design of RGB-D backbones for feature extraction has been mostly based on two or three stream architectures [8]. However, the performance of different frameworks depends largely on how they fuse RGB and Depth features. Some early works, such as FuseNet[36] and RedNet [37], fused RGB and Depth feature maps pixel-wise in the backbone. Later works, such as ACNet [25], ESANet [9] and EMSANet[11], proposed channel attention to select features from different channels, as RGB and Depth feature maps may not align well on the corresponding channels. PSCNet [12] further extended channel attention to both spatial and channel directions and achieved better performance. Recently, more complex models have been proposed to exploit cross-modal information and select features for RGB-D fusion. For example, SAGate [27] proposed a gated attention mechanism that can leverage cross-modal information for feature selection. CANet [28] extended non-local attention [20] to cross-modal semantic information and achieved significant improvement. CMX [8] extended SA-Gate to spatial and channel directions and proposed a novel cross-modal attention with global receptive field, which can enhance cross-modal complementarity using the similarity of two modalities. However, integrating cross-modal information and learning cross-modal similarity is still an open question in vision tasks.
III Method
Figure 1: Overview of AsymFormer.
III-A Framework Overview
This paper proposes AsymFormer, a novel network for RGB-D integration semantic segmentation. The network employs a ConvNext [16] based backbone to extract features from RGB images and a Mix-Transformer [17] based backbone to process RGB-D fused features. We introduced the LAFS module for feature selection, utilizing learnable feature weights to calculate spatial attention weights. We integrated modeling of self-similarity in multi-modal features, validating its capability to enhance network accuracy with minimal additional model parameters. The overall framework of AsymFormer is illustrated in Figure 1.
III-B Local Attention-Guided Feature Selection
Attention mechanisms have been shown to be effective in selecting complementary features from RGB and Depth features, thus improving efficiency and performance[25, 1, 12, 8]. Current spatial attention methods extract global information using pixel maximum or average values, neglecting the differences in channels from multiply modalities. In this work, we extend the previous spatial attention by proposing a more flexible spatial attention mechanism. The core design of LAFS is a trainable spatial attention weight. This weight is learned through an MLP in a similar manner to SE [34].
In specific, Figure 1 (b)-1 illustrates the details of LAFS. The LAFS receives input features of the concatenation of RGB feature $RGB$ with dimension $(C_{1},H,W)$ and current layer Depth feature $Depth$ with dimension $(C_{2},H,W)$.
For the computation of the attention weight, we first extract global information vectors $Avg$ of each feature map by adaptive AvgPool. For the channel attention, we process the global information by a feed-forward network with a squeeze-and-excitation structure.
$W_{C}=\text{Sigmoid}(\text{FFN}(Avg))$ (1)
For the spatial attention, we process the global information $Avg$ by another feed-forward network with a squeeze-and-excitation structure. The output vector is split into $R_{Max}$ and $R_{Avg}$, which represent two different descriptions of pixel spatial-similarity.
$R_{Avg}=\text{FFN}(Avg)$ (2)
Furthermore, the global spatial information $I_{S}$ of feature map can be extracted by calculating the inner product similarity from $R_{Avg}$ and the $Input$ (In computation, this is equivalent to the weighted sum of the features from each channel). Then, the spatial attention weight $W_{S}$ is computed by sigmoid normalization.
$W_{S}=\text{Sigmoid}(\frac{\text{Dot}(Input.\text{Reshape}(C,H\times W)^{T},R_{Avg})}{C^{2}})$ (3)
[…]
$Output=W_{C}\times W_{S}\times Input$ (4)
III-C Cross Modal Attention-Guided Feature Correlation Embedding
In this paper, we introduce a novel cross-modal attention (CMA) module that incorporates cross-modal self-similarity information into pixel-wise fused features. Current MHSA [21] is limited to learning the self-similarity of a single modality, whereas in multi-modal processing, our goal is to jointly use information from multiple modalities for representation learning. To achieve this, CMA is proposed to learn the self-similarity of multiple modalities. The key insight of CMA is independently learning the self-similarity of two modalities and embeds the sum of the results into the fusion feature, which is different than standard MHSA in theoretical. Specifically, for each pixel $(i,j)$ in feature map, we separately embed its RGB and Depth feature:
$Key_{RGB}=(Kr_{1,i,j},Kr_{2,i,j},...,Kr_{N,i,j})$ (5)
$Query_{RGB}=(Qr_{1,i,j},Qr_{2,i,j},...,Qr_{N,i,j})$
$Key_{Depth}=(Kd_{1,i,j},Kd_{2,i,j},...,Kd_{N,i,j})$
$Query_{Depth}=(Qd_{1,i,j},Qd_{2,i,j},...,Kd_{N,i,j})$
The cross-modal similarity of pixel $(i_{0},j_{0})$ and other pixel in feature map are defined as:
$W(i,j)=\sum_{n=1}^{N}(Kr_{n,i,j}\cdot Qr_{n,i_{0},j_{0}})+\sum_{n=1}^{N}(Kd_{n,i,j}\cdot Qd_{n,i_{0},j_{0}})\$ (6)
Moreover, the CMA embeds cross-modal similarity into the fused feature $Value$ with same operation as[21, 20, 22]. Following above idea, the CMA first embedding $Key$, $Query$ and $Value$ independently for three modalities features. Then, we extend self-attention to the multimodal case, which contain two design: 1. Information mixing and redistribution of $Key$ and $Query$; 2. Learning two different different representation subspaces for the same position.
Independent linear embedding: According to our definition, the $Key$, $Query$ and $Value$ should be embedded independently. Assuming the input $RGB$ feature has dimension ${}^{(}C_{0},H,W)$; the Enhanced Depth feature $Depth_{En.}$ has dimension ${}^{(}C_{1},H,W)$; and the previous fused feature $Fus_{1}$ has dimension $(C_{2},H,W)$. We first normalize the RGB feature and Depth feature. Then, we use six independent convolution layers to generate self-attention keys and queries from RGB and Depth feature:
$RGB_{emb}=\left\{\begin{array}[]{c}K_{RGB}=\text{Conv}_{f\times f}(RGB)\\ Q_{RGB}=\text{Conv}_{1\times 1}(RGB)\\ \end{array}\right.$ (7)
$Depth_{emb}=\left\{\begin{array}[]{c}K_{Depth}=\text{Conv}_{f\times f}(Depth_{En.})\\ Q_{Depth}=\text{Conv}_{1\times 1}(Depth_{En.})\\ \end{array}\right.$ (8)
$Fused_{emb}=\left\{\begin{array}[]{c}V=\text{Conv}_{f\times f}(Fused_{1})\\ V_{1},V2=\text{Split}(V,2)\\ \end{array}\right.$ (9)
The dimensions of $K_{RGB}$ and $K_{Depth}$ are $(\frac{C_{1}}{4},\frac{H}{f},\frac{W}{f})$, where $f$ is the down-sample rate and equals the convolution kernel-size and stride. This reduces the feature map resolution to $\frac{1}{f}$. The dimensions of $Q_{RGB}$ and $Q_{Depth}$ are $(\frac{C_{1}}{4},H,W)$. The $V_{1}$ and $V_{2}$ is reshaped into dimension $(\frac{C_{2}}{4},\frac{H\times W}{f^{2}})$.
Information mixing and redistribution: We concatenate the cross-modal keys and queries to obtain the $Key$ and $Query$:
$Key,Query=\left\{\begin{array}[]{c}Key=\text{Cat}[K_{RGB},K_{Depth}]\\ Query=\text{Cat}[Q_{RGB},Q_{Depth}]\\ \end{array}\right.$ (10)
The first $\frac{C_{2}}{4}$ channels of $Key$ and $Query$ contain RGB information. Similarly, the second half of the $Key$ and $Query$ channels contain Depth information. To obtain the $K1$, $K2$ and $Q1$, $Q2$ for integrating RGB-D information computing the two feature subspaces $W_{1}$, $W_{2}$, we perform a channel shuffle of $Key$ and $Query$ and split them into required vector:
$K_{1},K_{2}=\text{Split}(\text{Shuffle}(Key),2)$ (11)
$Q_{1},Q_{2}=\text{Split}(\text{Shuffle}(Query),2)$
Representation subspaces learning: The $Q_{1}$, $Q_{2}$ are reshaped into dimension $(\frac{C_{1}}{4},H\times W)$ and the $K_{1}$, $k_{2}$ are reshaped into dimension $(\frac{C_{1}}{4},\frac{H\times W}{f^{2}})$. Then, the two representation subspaces $W_{1}$ and $W_{2}$ are computed by dot product of $K$ and $Q$ respectively:
$W_{1}=\text{Softmax}(\frac{Q_{1}\cdot K_{1}^{T}}{\sqrt{C_{1}/4}})$ (12)
$W_{2}=\text{Softmax}(\frac{Q_{2}\cdot K_{2}^{T}}{\sqrt{C_{1}/4}})$
The $W_{1}$ and $W_{2}$ are embedded into $V_{1}$ and $V_{2}$ by dot product. The fused feature $Fused_{2}$ is the concatenation of $V_{1}$ and $V_{2}$:
$Fused_{2}=\text{Cat}[W_{1}\cdot V_{1},W_{2}\cdot V_{2}]\$ (13)
Finally, the $Fused_{2}$ is adapted into output channel $C_{2}$ and added with residual connection $Fused_{1}$.
$Fused_{2}=\text{ConvBN}_{1\times 1}(Fused_{2})+Fuesed_{1}$ (14)
TABLE I: Ablation experiment results for different multi-modal feature fusion method.
Model Feature Fusion Method Metric
Cat SE+MHSA LAFS CMA Params/M mIoU(%) Inf.Speed(FPS)
Baseline 31.9 47.0 77.5
32.6(+0.7M) 49.9 (+2.9%) 65.7
Ours 32.4 (+0.5M) 49.1 (+2.1%) 75.7
32.5 (+0.6M) 49.6 (+2.6%) 67.4
33.0 (+1.1M) 54.1 (+7.1%) 65.5
IV EXPERIMENT RESULTS AND ANALYSIS
IV-A Implementation Details
To evaluate our Real-Time semantic segmentation network design, we conduct a series of experiments on two widely-used datasets: NYUv2[5] (795 training and 654 testing RGB-D images) and SUNRGBD[6] (5825 training and 5050 testing RGB-D images).
We conduct the model training and testing on different platforms. For the training, we use Nvidia A100-40G GPU. For the evaluation and inference speed testing, we use Nvidia RTX 3090 GPU, Ubuntu 20.04, CUDA 12.0 and Pytorch 2.0.1. We apply data augmentation to all datasets by randomly flipping (p=0.5), srandom scales [1.0,2.0], random crop 480$\times$640 and random HSV. We adopt Mix Transformer[17] and ConvNext[16] backbones pretrained on ImageNet-1k[24]. In the augmented setting, both backbones are pretrained on ADE20k[23]. For the CMA implementation, we set down-sample rate $f$=1 in CMA-1 and $f$=2 in CMA-2,3. The MLP-decoder in AsymFormer has the same structure as Segformer and an embedding dimension of 256. We choose AdamW optimizer with a weight decay of 0.01. The initial learning rate is $5e^{-5}$ and we use a poly learning rate schedule $(1-\frac{iter}{max_{i}ter})^{0.9r}$ with a warm-up of 10 epochs. We train with a batch size of 8 for NYUv2 (500 epochs) and SUNRGBD (200 epochs). We employ cross-entropy as the loss function and do not use any auxiliary loss during training process. The evaluation metric is mean Intersection over Union (mIoU).
IV-B Model Assessment
This paper proposes two different models based on the AsymFormer framework with different backbones. The AsymFormer uses ConvNext-T[16] for RGB representation learning and MiT-B0[17] for fused feature processing. This model has 33.0 million parameters and 36.0 GFLOPs computational cost, and it can achieve 65 FPS inference speed on RTX 3090. Particularly, with mixed precision quantization, AsymFormer achieves a real-time inference speed of 79 FPS on the RTX 3090 GPU. It also achieves a real-time inference speed of 29.4 FPS on the Tesla T4 GPU, which has a more limited computation resources (65 TFLOPs FP16).
IV-C Ablation Experiment
We conduct a series of ablation experiments on NYUv2 dataset to evaluate the effectiveness of the LAFS and CMA module. We set two common feature fusion methods as our comparative baseline: 1. Cat: This method directly concatenates two features and then uses convolution layers to adjust the channel numbers. Essentially, it is a pixel-wise fusion without feature selection. 2. SE+MHSA: This method combines the popular SE attention [34] and MHSA attention [21] for feature fusion. Here, SE is used for feature selection in the channel direction, while MHSA is employed for further feature extraction on the fused features. Specifically, we use an MHSA with 2 heads to align with CMA.
In our experiments, the Cat fusion method, used as a baseline, achieved a segmentation accuracy of 47.0 mIoU and an inference speed of 77.5 FPS. When using LAFS for feature selection alone, we achieved a performance improvement of 2.1% while sacrificing only 1.8 FPS of inference speed. This demonstrates that LAFS is a highly efficient feature selection method that provides performance gains without significantly impacting inference speed. In comparison to the other baseline, Cat+MHSA, which resulted in an 11.8 FPS reduction in inference speed, only a 2.9% improvement in segmentation accuracy was achieved. This further highlights the efficiency of LAFS.
Furthermore, we conducted experiments to investigate whether incorporating the mining of self-similarity within the two modalities after feature selection could further improve segmentation accuracy. When using CMA alone, we observed a 2.6% improvement in segmentation accuracy but encountered a significant decrease in inference speed of 10.1 FPS. Compared to LAFS, CMA showed a more noticeable reduction in inference speed. In comparison to SE+MHSA, CMA achieved similar accuracy, but it is important to note that no feature selection was performed in this case, indicating that relying solely on mining multi-modal self-similarity can achieve accuracy similar to the traditional feature fusion baseline SE+MHSA.
Finally, we combined LAFS with CMA (LAFS+CMA). Since LAFS had minimal impact on inference speed and served a different purpose than CMA, the network’s inference speed only decreased by 2 FPS while achieving a significant improvement of 7.1% compared to the baseline Cat. At this point, the inference speed of LAFS+CMA was similar to SE+MHSA, but with a 2.1% performance improvement. This validates our experimental hypothesis: by re-modeling feature selection and mining cross-modal self-similarity, we can enhance the segmentation performance of the network without sacrificing inference speed compared to existing models. This demonstrates that we have indeed improved the efficiency of the network.
IV-D Comparison With State-of-The-Arts
IV-D1 NYUv2 Comparison Results
[…]
TABLE II: Comparison Results on NYUv2. * denotes ADE20k pretrain. The inference speed is tested on RTX 3090 platform, (480 $\times$ 640) inputs.
Method Year Backbone Params/M Pixel Acc. (%) mIoU (%) Inf Speed/FPS Inf Speed (%)
PSCNet-T 2022 Res50 40 - 45.4 47 72.3%
PSCNet-L 2022 Res50 52 - 46.2 30 46.2%
ACNet 2019 Res50 118 - 48.3 28 43.1%
RDFNet 2017 Res152 450 76 49.1 15 23.1%
SA-Gate 2021 Res50 65 76.8 50.2 35 53.8%
ESANet 2022 Res34-Nbt1D 34 - 51.6 32 49.2%
MSFNet 2022 Res50 49 - 52.2 25 38.5%
FRNet 2022 Res34 86 77.6 53.6 39 60.0%
PGDENet 2022 Res34 101 78.1 53.7 32 49.2%
CMX-B2 2022 Segformer-B2 67 78.7 54.1 24 36.9%
AsymFormer 2023 B0+T 33 78.5 54.1 65 100.0%
AsymFormer (Multi-Scale) 2023 B0+T 33 79.3 55.3 - -
AsymFormer (FP16) 2023 B0+T 33 78.5 54.1 79 121.5%
IV-D2 SUNRGBD Comparison Results
Table III reports the performance of AsymFormer on the SUNRGBD dataset. AsymFormer achieves competitive accuracy with 49.1 % mIoU. Since the input image size is the same as NYUv2, all methods have the same inference speed as shown in Table II. However, the advantage of AsymFormer is not as significant as in NYUv2 experiment. For example, AsymFormer improves 3.9 mIoU over SA-Gate[27] in NYUv2 dataset (54.1% vs 50.2% mIoU), but decreases 0.3 mIoU in SUNRGBD dataset (49.1% vs 49.4% mIoU). A similar performance degradation can be observed in CMX-B2 result which also uses Transformer based backbone. We conjecture that this phenomenon may be caused by low quality depth images in SUNRGBD dataset. The aim of our research is not to construct a state-of-the-art method that has a marginal mIoU improvement over other methods, but to construct a method that has a better performance-speed balance and is more suitable for robot platform. Given that AsymFormer still has faster inference speed than other methods, we consider this performance acceptable for AsymFormer.
TABLE III: Comparison Results on SUNRGBD. # denotes multi-scale test.
Method Pixel Acc. (%) mIoU (%)
RDFNet[26] 81.5 47.7
ESANet[9] - 48.0
ACNet[25] - 48.1
SA-Gate[27] 82.5 49.4
CMX-B2[8] 82.8 49.7
MSFNet[30] - 50.3
FRNet[10] 87.4 51.8
PGDENet[29] 87.7 51.0
AsymFormer 81.9 49.1
IV-E Visualization
1) LAFS Attention Map: As shown in Figure 2, to demonstrate that LAFS performs better than CBAM [18] in selecting features in the spatial dimension, we visualized the spatial attention weights of both methods. It can be observed that LAFS provides better coverage of informative regions in the image while maintaining consistency within objects and preserving the integrity of edges.
Figure 2: Difference between CBAM and LAFS’s spatial attention map.
2) Semantic Segmentation Results: The Figure 3 demonstrates the segmentation results of AsymFormer on the NYUv2 dataset. As observed, while maintaining a significantly faster inference speed compared to other methods, AsymFormer achieves comparable semantic segmentation accuracy to mainstream approaches.
Figure 3: Visualization of AsymFormer Semantic Segmentation Results.
V CONCLUSIONS
In this work, we proposed AsymFormer, which aims to construct a less-redundant real-time indoor scene understanding system. To enhance efficiency and reduce redundant parameters, we implemented the following improvement: 1. We employed an asymmetric backbone that compressed the parameters of the Depth feature extraction branch, thus reducing redundancy. 2. We introduced the LAFS module for feature selection, utilizing learnable feature weights to calculate spatial attention weights. This resulted in a significant performance improvement while almost maintaining the inference speed. 3. Expanding on feature selection, we integrated modeling of self-similarity in multi-modal features, validating its capability to enhance network accuracy with minimal additional model parameters. The experiments demonstrated that the AsymFormer surpassed existing SOTA methods in terms of speed, demonstrating our method achieved a balance between accuracy and speed. Moving forward, we will continue to optimize the modules and address issues such as self-supervised pre-training of the model, aiming for further improvements.

### v7

@@
 Abstract
-In the realm of robotic intelligence, achieving efficient and precise RGB-D semantic segmentation is a key cornerstone.
-State-of-the-art multimodal semantic segmentation methods, primarily rooted in symmetrical skeleton networks, find it challenging to harmonize computational efficiency and precision.
-In this work, we propose AsymFormer, a novel network for real-time RGB-D semantic segmentation, which targets the minimization of superfluous parameters by optimizing the distribution of computational resources and introduces an asymmetrical backbone to allow for the effective fusion of multimodal features.
-Furthermore, we explore techniques to bolster network accuracy by redefining feature selection and extracting multi-modal self-similarity features without a substantial increase in the parameter count, thereby ensuring real-time execution on robotic platforms.
-Additionally, a Local Attention-Guided Feature Selection (LAFS) module is used to selectively fuse features from different modalities by leveraging their dependencies.
+Understanding indoor scenes is crucial for urban studies.
+Considering the dynamic nature of indoor environments, effective semantic segmentation requires both real-time operation and high accuracy.To address this, we propose AsymFormer, a novel network that improves real-time semantic segmentation accuracy using RGB-D multi-modal information without substantially increasing network complexity.
+AsymFormer uses an asymmetrical backbone for multimodal feature extraction, reducing redundant parameters by optimizing computational resource distribution.
+To fuse asymmetric multimodal features, a Local Attention-Guided Feature Selection (LAFS) module is used to selectively fuse features from different modalities by leveraging their dependencies.
 Subsequently, a Cross-Modal Attention-Guided Feature Correlation Embedding (CMA) module is introduced to further extract cross-modal representations.
-This method is evaluated on NYUv2 and SUNRGBD datasets, with AsymFormer demonstrating competitive results with 54.1% mIoU on NYUv2 and 49.1% mIoU on SUNRGBD.
-Notably, AsymFormer achieves an inference speed of 65 FPS and after implementing mixed precision quantization, it attains an impressive inference speed of 79 FPS on RTX3090.
-This significantly outperforms existing multi-modal methods, thereby demonstrating that AsymFormer can strike a balance between high accuracy and efficiency for RGB-D semantic segmentation.
+The AsymFormer demonstrates competitive results with 54.1% mIoU on NYUv2 and 49.1% mIoU on SUNRGBD.
+Notably, AsymFormer achieves an inference speed of 65 FPS (79 FPS after implementing mixed precision quantization) on RTX3090, demonstrating that AsymFormer can strike a balance between high accuracy and efficiency.
 Code: https://github.com/Fourier7754/AsymFormer
-I INTRODUCTION
-Real-time semantic segmentation methods are important for robots as they rely on obtaining semantic information in real-time to support navigation and task decision-making.
-Current Real-Time semantic segmentation networks achieve competitive results and reach over 100 FPS inference speed on simple outdoor environments (Cityscapes 19 classes) [3, 4, 2, 1].
-However, these networks perform poorly on complex indoor environments, such as NYU Depth v2 (40 classes) [5] and SUNRGBD (37 classes) [6].
-Several studies have explored how to improve indoor scene semantic segmentation performance by integrating RGB-D information [7, 8, 9, 10], which is widely available from RGB-D sensors on indoor robots.
-RGB-D information consists of RGB (color, texture and shape) and Depth (boundaries and relative location) features, which are somewhat complementary [7, 8].
-At present, RGB-D semantic segmentation methods have achieved state-of-the-art performance by employing dual-branch backbone for feature extraction and attention mechanisms for feature selection [8].
-However, there is a dearth of discourse on whether these operations can be executed with less redundancy and higher efficiency.
-Primarily, most methodologies utilize two symmetrical, computationally intensive backbones, which effectively doubles the overall computational complexity [12].
-Secondly, numerous methods concentrate on designing complex feature fusion modules to enhance accuracy, but do not exhibit significant advantages than some methods that employ simpler attention modules like ESANet [9], which indicates that as more computational resources are invested, the incremental improvement in accuracy decreases.
-Lastly, the use of cross-attention modules to jointly utilize multi-modal features has been demonstrated to be effective in improving network accuracy [8, 28].
-However, most methods do not discuss how to implement cross-attention more efficiently.
-Therefore, the reduction of redundancy in dual-branch networks and the construction of more efficient attention modules remain significant topics for discussion.
-To address this issue, this study suggests a balanced approach between computational efficiency and feature richness in multi-modal representation learning.
-RGB images are found to contain more information than Depth images, thus it is not necessary to use the same size backbone for extracting Depth information.
-The performance improvements achieved by using attention mechanisms for various tasks come with minimal increase in computational cost.
-While existing modules like SE [34] are efficient for feature selection, they lack the ability to model correlations in spatial dimensions.
-Increasing parameters can facilitate spatial dimension feature selection, but careful design is needed to maintain computational efficiency.
-For feature extraction, the commonly used Multi-Head Self Attention (MHSA) [22] is effective for single-modal features despite its small parameter count.
-Extending this efficient feature extraction to multiple modalities is feasible, but consideration should be given to feature fusion while maintaining computational efficiency.
-This paper introduces AsymFormer, a high-performance real-time network for RGB-D integration semantic segmentation that employs an asymmetrical backbone design.
-This includes a larger parameter backbone for RGB features and a smaller backbone for the Depth branch.
-Regarding framework selection, Vision Transformer are always performance better, but slower than same computational complexity CNNs due to a lack of hardware optimization [40].
-In order to speed up the main branch, this paper employ a hardware-friendly CNN [16] for the RGB branch and a Transformer [17] with fewer parameters but similar performance for the Depth branch to further compress the parameters.
-This paper also discuss feature selection modeling and construct the Local Attention-Guided Feature Selection (LASF) module.
-This module computes the spatial-channel attention weights in parallel to enhance inference speed.
-Moreover, this paper establishes a novel approach to model spatial attention.
-This module can estimate the differences in features from different modalities by utilizing learnable channel weights and calculating spatial attention weights for each pixel.
-Additionally, a Cross-Modal Attention (CMA) module is introduced to embed cross-modal integration learning information into pixel-wise fused features.
+111$\nmid$ indicates equal contribution.222* corresponding author.
+Figure 1: The AsymFormer has 33.0 million parameters and 36.0 GFLOPs computational cost, and it can achieve 65 FPS inference speed on RTX 3090, 54.1% mIoU on NYUv2.
+1 Introduction
+Indoor scenes are essential to urban environments.
+Current urban studies necessitate understanding indoor scene semantic information for tasks like emergency evacuation [38], robotic navigation [36], and virtual reality [43].
+The dynamic nature of indoor scenes demands perception of environmental changes in real time for tasks such as emergency evacuation [38], stressing the importance of algorithms’ real-time capabilities.
+Existing real-time semantic segmentation methods often falter when applied to the complex semantic information of indoor scenes, usually requiring a sacrifice in accuracy [40].
+This often requires a trade-off between inference speed and increased network complexity to achieve adequate segmentation accuracy indoors.
+Consequently, a critical research question is how to improve the accuracy of semantic segmentation in indoor environments without substantially increasing complexity, while ensuring real-time performance.
+Aside from increasing the complexity of the network, introducing additional information, such as RGB-D data, is also an effective way to improve the accuracy of semantic segmentation networks.
+RGB-D cameras are widely used devices for indoor information acquisition.
+RGB-D information consists of RGB (color, texture and shape) and Depth (boundaries and relative location) features, which are somewhat complementary [42].
+Several studies have explored how to improve indoor scene semantic segmentation performance by integrating RGB-D information [42, 26, 50].
+Existing research has explored the implementation of attention mechanisms to extract valuable information from RGB-D features without significantly increasing computational complexity.
+However, due to the additional feature extraction branch for depth features and the lack of discussion on how to allocate computational resources based on feature importance, these methods often introduce a substantial amount of redundant parameters, significantly reducing their computational efficiency [9].
+To address this issue, this paper introduces AsymFormer, a high-performance real-time network for RGB-D semantic segmentation that employs an asymmetric backbone design.
+This includes a larger parameter backbone for important RGB features and a smaller backbone for the Depth branch.
+Regarding framework selection, at the same computational complexity, Transformer often achieves higher accuracy but has a slower inference speed compared to CNN [19].
+In order to speed up the main branch, this paper employ a hardware-friendly CNN [21] for the RGB branch and a light-weight Transformer [35] for the Depth branch to further compress the parameters.
+Considering the differences between different modal representation, to effectively select and fuse asymmetric features, this paper proposes a learnable method for feature information compression and constructs a Local Attention Guided Feature Selection (LAFS) module.
+Additionally, a Cross-Modal Attention (CMA) module is introduced to embed cross-modal information into pixel-wise fused features.
 Finally, we employ a lightweight MLP-Decoder[35] to decode semantic information from shallow features.
@@
 Meanwhile, the inference speed test is also preformed on Nvidia RTX 3090 platform.
-The AsymFormer achieves 54.1% mIoU on NYUv2 and 49.1 mIoU on SUNRGBD, with 65 FPS inference speed.
-After performing mixed precision quantization on the model using the TensorRT framework, AsymFormer achieves an impressive inference speed of 79 FPS on RTX 3090, which significantly surpasses existing methods.
+The AsymFormer achieves 54.1% mIoU on NYUv2 and 49.1 mIoU on SUNRGBD, with 65 FPS inference speed (79 FPS with mixed precision quantization using the TensorRT).
 Our experiments highlight AsymFormer’s ability to acquire high accuracy and efficiency at the same time.
-Our main contributions are summarized as follows:
+The main contributions are summarized as follows:
   • $\bullet$
@@
     We introduce a novel efficient cross-modal attention (CMA) for modeling of self-similarity in multi-modal features, validating its capability to enhance network accuracy with minimal additional model parameters.
-II Related Works
-RGB-D Representation Learning.
+2 Related Works
+Figure 2: Overview of AsymFormer.
+2.1 Indoor Scene Understanding
+Current research in indoor scene understanding leverages diverse data sources, including RGB images [2], RGB-D images [42], point clouds [51], and mesh data [45].
+Each of these sources provides unique insights and benefits for analyzing and interpreting indoor environments.
+For instance, RGB images are accessible and straightforward for visual representations, whereas point clouds and meshes offer intricate 3D spatial data.
+However, when considering the simultaneous use of both 2D and 3D information to optimize efficiency and effectiveness in scene understanding, RGB-D images emerge as the optimal choice.
+2.2 RGB-D Representation Learning
 One of the earliest works on RGB-D semantic segmentation, FCN [22], treated RGB-D information as a single input and processed it with a single backbone.
 However, subsequent works have recognized the need to extract features from RGB and Depth information separately, as they have different properties.
-Therefore, most of them have adopted two symmetric backbones for RGB and Depth feature extraction [36, 37, 9, 11, 27, 8].
-Further, the ACNet [25] proposed a three-stream backbone to process RGB, Depth and Fusion features independently.
-Moreover, some works have considered the variability of RGB-D features and used asymmetric backbones for their feature extraction.
-For example, TSNet [39] used ResNet [31] for RGB feature extraction and VGG (without residual connections) [38] for Depth feature extraction.
-The PSCNet [12] also used an asymmetric backbone and reduced the redundant computational cost by cutting the parameters of the Depth branch.
-RGB-D feature fusion.
-The design of RGB-D backbones for feature extraction has been mostly based on two or three stream architectures [8].
-However, the performance of different frameworks depends largely on how they fuse RGB and Depth features.
-Some early works, such as FuseNet[36] and RedNet [37], fused RGB and Depth feature maps pixel-wise in the backbone.
-Later works, such as ACNet [25], ESANet [9] and EMSANet[11], proposed channel attention to select features from different channels, as RGB and Depth feature maps may not align well on the corresponding channels.
+Therefore, most of them have adopted two symmetric backbones for RGB and Depth feature extraction [26, 27, 5, 42].
+Primarily, asymmetric backbones doubles the overall computational complexity [9].
+However, it is generally observed that for semantic segmentation, RGB information typically plays a more prominent role than Depth information, as indicated by [9].
+Using a asymmetric backbone for feature extraction will obviously lead to redundant parameters on the less important side, reducing the efficiency of the network.
+2.3 RGB-D feature fusion
+The performance and efficiency of different frameworks depends largely on how they fuse RGB and Depth features.
+Some early works, such as RedNet [17], fused RGB and Depth feature maps pixel-wise in the backbone.
+Later, ESANet series [26, 27] proposed channel attention to select features from different channels, as RGB and Depth feature maps may not align well on the corresponding channels.
 PSCNet [9] further extended channel attention to both spatial and channel directions and achieved better performance.
@@
 CANet [47] extended non-local attention [32] to cross-modal semantic information and achieved significant improvement.
-CMX [8] extended SA-Gate to spatial and channel directions and proposed a novel cross-modal attention with global receptive field, which can enhance cross-modal complementarity using the similarity of two modalities.
+CMX [42] extended SA-Gate to spatial and channel directions and proposed a novel cross-modal attention with global receptive field.
 However, integrating cross-modal information and learning cross-modal similarity is still an open question in vision tasks.
-III Method
-Figure 1: Overview of AsymFormer.
-III-A Framework Overview
-This paper proposes AsymFormer, a novel network for RGB-D integration semantic segmentation.
-The network employs a ConvNext [16] based backbone to extract features from RGB images and a Mix-Transformer [17] based backbone to process RGB-D fused features.
-We introduced the LAFS module for feature selection, utilizing learnable feature weights to calculate spatial attention weights.
-We integrated modeling of self-similarity in multi-modal features, validating its capability to enhance network accuracy with minimal additional model parameters.
-The overall framework of AsymFormer is illustrated in Figure 1.
-III-B Local Attention-Guided Feature Selection
-Attention mechanisms have been shown to be effective in selecting complementary features from RGB and Depth features, thus improving efficiency and performance[25, 1, 12, 8].
-Current spatial attention methods extract global information using pixel maximum or average values, neglecting the differences in channels from multiply modalities.
-In this work, we extend the previous spatial attention by proposing a more flexible spatial attention mechanism.
-The core design of LAFS is a trainable spatial attention weight.
-This weight is learned through an MLP in a similar manner to SE [34].
-In specific, Figure 1 (b)-1 illustrates the details of LAFS.
-The LAFS receives input features of the concatenation of RGB feature $RGB$ with dimension $(C_{1},H,W)$ and current layer Depth feature $Depth$ with dimension $(C_{2},H,W)$.
-For the computation of the attention weight, we first extract global information vectors $Avg$ of each feature map by adaptive AvgPool.
-For the channel attention, we process the global information by a feed-forward network with a squeeze-and-excitation structure.
-$W_{C}=\text{Sigmoid}(\text{FFN}(Avg))$ (1)
-For the spatial attention, we process the global information $Avg$ by another feed-forward network with a squeeze-and-excitation structure.
-The output vector is split into $R_{Max}$ and $R_{Avg}$, which represent two different descriptions of pixel spatial-similarity.
-$R_{Avg}=\text{FFN}(Avg)$ (2)
-Furthermore, the global spatial information $I_{S}$ of feature map can be extracted by calculating the inner product similarity from $R_{Avg}$ and the $Input$ (In computation, this is equivalent to the weighted sum of the features from each channel).
-Then, the spatial attention weight $W_{S}$ is computed by sigmoid normalization.
+3 Method
+3.1 Framework Overview
+This paper develops a high-accuracy real-time semantic segmentation method, AsymFormer, which leverages an asymmetric backbone network to mine multimodal representations.
+A hardware friendly convolution network ConvNext [21] is used for RGB feature extraction and a light-weight Mix-Transformer [35]is used for processing Depth feature.
+To effectively fuse RGB-D features, the study introduces a Local Attention Guided Multimodal Feature Selection (LAFS) module, which uses learnable strategy to extract global information and selects multimodal features in both spatial and channel dimensions.
+Moreover, to further mine the information contained in multimodal features, the study embeds the information contained in multimodal features through a novel Cross-Modal Attention (CMA) module.
+The overall framework of AsymFormer is shown in Fig.2.
+3.2 Local Attention Guided Feature Selection
+Studies have shown that attention mechanisms can effectively select complementary features from RGB and Depth features, thereby improving the efficiency of feature extraction and the performance of semantic segmentation [16, 40, 9, 42].
+However, existing attention mechanisms usually adopt a non-learnable, fixed strategy for feature information compression [34], which may overlook the differences among various modal features, causing insufficient use of information.
+Figure 3: LAFS.
+In response to this issue, this paper proposes a learnable method for feature information compression and constructs a Local Attention Guided Feature Selection (LAFS) module.
+The core design of LAFS is a learnable method for spatial information compression.
+Inspired by SE [14] attention, LAFS abandons traditional fixed strategies and employs a feedforward neural network to learn a set of dynamic spatial information compression rules.
+The Fig.
+3 demonstrates details of the LAFS module.
+The input features to the LAFS module are the tensor concatenation of $RGB$ and $Depth$ features.
+For computing attention weights, the study first extracts a global information vector $Avg$ through Adaptive Average Pooling.
+For channel attention, its structure is the same as SE [14].
+On spatial attention computation, which is the primary improvement, the LAFS first processes global information $Avg$ through another feedforward neural network with a squeeze-excitation structure, producing the output vector $R_{Avg}$, representing pixel space similarity descriptions.
+Then, by computing the dot product similarity between $R_{Avg}$ and input feature map, the global spatial information $I_{S}$ is extracted (this equates to a weighted sum of the channel features, dynamically varying with $Avg$ changes).
+Subsequently, spatial attention weights $W_{S}$ are calculated through sigmoid normalization.
 $W_{S}=\text{Sigmoid}(\frac{\text{Dot}(Input.\text{Reshape}(C,H\times W)^{T},R_{Avg})}{C^{2}})$ (1)
-[…]
-$Output=W_{C}\times W_{S}\times Input$ (4)
-III-C Cross Modal Attention-Guided Feature Correlation Embedding
-In this paper, we introduce a novel cross-modal attention (CMA) module that incorporates cross-modal self-similarity information into pixel-wise fused features.
-Current MHSA [21] is limited to learning the self-similarity of a single modality, whereas in multi-modal processing, our goal is to jointly use information from multiple modalities for representation learning.
-To achieve this, CMA is proposed to learn the self-similarity of multiple modalities.
-The key insight of CMA is independently learning the self-similarity of two modalities and embeds the sum of the results into the fusion feature, which is different than standard MHSA in theoretical.
-Specifically, for each pixel $(i,j)$ in feature map, we separately embed its RGB and Depth feature:
-$Key_{RGB}=(Kr_{1,i,j},Kr_{2,i,j},...,Kr_{N,i,j})$ (5)
-$Query_{RGB}=(Qr_{1,i,j},Qr_{2,i,j},...,Qr_{N,i,j})$
-$Key_{Depth}=(Kd_{1,i,j},Kd_{2,i,j},...,Kd_{N,i,j})$
-$Query_{Depth}=(Qd_{1,i,j},Qd_{2,i,j},...,Kd_{N,i,j})$
-The cross-modal similarity of pixel $(i_{0},j_{0})$ and other pixel in feature map are defined as:
-$W(i,j)=\sum_{n=1}^{N}(Kr_{n,i,j}\cdot Qr_{n,i_{0},j_{0}})+\sum_{n=1}^{N}(Kd_{n,i,j}\cdot Qd_{n,i_{0},j_{0}})\$ (6)
-Moreover, the CMA embeds cross-modal similarity into the fused feature $Value$ with same operation as[21, 20, 22].
-Following above idea, the CMA first embedding $Key$, $Query$ and $Value$ independently for three modalities features.
-Then, we extend self-attention to the multimodal case, which contain two design: 1.
-Information mixing and redistribution of $Key$ and $Query$; 2.
-Learning two different different representation subspaces for the same position.
-Independent linear embedding: According to our definition, the $Key$, $Query$ and $Value$ should be embedded independently.
-Assuming the input $RGB$ feature has dimension ${}^{(}C_{0},H,W)$; the Enhanced Depth feature $Depth_{En.}$ has dimension ${}^{(}C_{1},H,W)$; and the previous fused feature $Fus_{1}$ has dimension $(C_{2},H,W)$.
-We first normalize the RGB feature and Depth feature.
-Then, we use six independent convolution layers to generate self-attention keys and queries from RGB and Depth feature:
-$RGB_{emb}=\left\{\begin{array}[]{c}K_{RGB}=\text{Conv}_{f\times f}(RGB)\\ Q_{RGB}=\text{Conv}_{1\times 1}(RGB)\\ \end{array}\right.$ (7)
-$Depth_{emb}=\left\{\begin{array}[]{c}K_{Depth}=\text{Conv}_{f\times f}(Depth_{En.})\\ Q_{Depth}=\text{Conv}_{1\times 1}(Depth_{En.})\\ \end{array}\right.$ (8)
-$Fused_{emb}=\left\{\begin{array}[]{c}V=\text{Conv}_{f\times f}(Fused_{1})\\ V_{1},V2=\text{Split}(V,2)\\ \end{array}\right.$ (9)
-The dimensions of $K_{RGB}$ and $K_{Depth}$ are $(\frac{C_{1}}{4},\frac{H}{f},\frac{W}{f})$, where $f$ is the down-sample rate and equals the convolution kernel-size and stride.
-This reduces the feature map resolution to $\frac{1}{f}$.
-The dimensions of $Q_{RGB}$ and $Q_{Depth}$ are $(\frac{C_{1}}{4},H,W)$.
-The $V_{1}$ and $V_{2}$ is reshaped into dimension $(\frac{C_{2}}{4},\frac{H\times W}{f^{2}})$.
-Information mixing and redistribution: We concatenate the cross-modal keys and queries to obtain the $Key$ and $Query$:
+Here, all results are divided by $C^{2}$ to prevent sigmoid overflow.
+Finally, input features are selected based on the pixel-wise multiplication of feature maps with spatial and channel attention weights.
+3.3 Cross-Attention Guided Feature Embedding
+Aside from selecting existing features, it is necessary to mine new information from fused features.
+Existing Multi-Head Self-Attention (MHSA) in Transformer [31] is limited to learning self-similarity within a single modality, whereas mining information from multi-modal jointly is a new goal in representation learning.
+In this paper, a new Cross-Modal Attention (CMA) module is constructed.
+The key to CMA is defining cross-modal self-similarity using a linear sum, and embedding its result into the fused features.
+3.3.1 Definition of Cross-Modal Self-Similarity:
+Suppose $RGB$ and $Depth$ features are seamlessly embedded into $Key$ and $Query$, then for a pixel $(i_{0},j_{0})$, its cross-modal self-similarity with other pixels $(i,j)$ can be defined as:
+$W(i,j)=\sum_{n=1}^{N}(Kr_{n,i,j}\cdot Qr_{n,i_{0},j_{0}})+\sum_{n=1}^{N}(Kd_{n,i,j}\cdot Qd_{n,i_{0},j_{0}})$ (2)
+where $Kr_{n,i,j}$ represents the nth feature value of the pixel $Key_{RGB}$, and $Qr_{n,i,j}$ represents the nth feature value of the pixel $Query_{RGB}$.
+Similarly, $Kd_{1,i,j}$ and $Qd_{1,i,j}$ represent the nth feature value of the pixel $Key_{Depth}$ and $Query_{Depth}$ respectively.
+3.3.2 Feature Embedding
+The CMA has three input features, namely $RGB$ features, $Depth$ features, and the fused features $Fused$ selected by LAFS.
+In the calculation process of CMA, the first step is to embed the input features $RGB$, $Depth$, and $Fused$ into vector space.
+Figure 4: Feature Embedding.
+Simultaneously, the output $Value$ split into two independent vectors $V_{1}$ and $V_{2}$ in the channel direction.
+3.3.3 Splitting and Mixing of Multimodal Information
+To learn features from multiple subspaces, the embedded features $K_{RGB}$, $K_{Depth}$, $Q_{RGB}$, and $Q_{Depth}$ are split into two independent vectors.The study first concatenates $Key_{RGB+Depth}$ with $Query_{RGB+Depth}$ in the channel direction, obtaining $Key$ and $Query$:
 $Key,Query=\left\{\begin{array}[]{c}Key=\text{Cat}[K_{RGB},K_{Depth}]\\ Query=\text{Cat}[Q_{RGB},Q_{Depth}]\\ \end{array}\right.$ (3)
-The first $\frac{C_{2}}{4}$ channels of $Key$ and $Query$ contain RGB information.
-Similarly, the second half of the $Key$ and $Query$ channels contain Depth information.
-To obtain the $K1$, $K2$ and $Q1$, $Q2$ for integrating RGB-D information computing the two feature subspaces $W_{1}$, $W_{2}$, we perform a channel shuffle of $Key$ and $Query$ and split them into required vector:
-$K_{1},K_{2}=\text{Split}(\text{Shuffle}(Key),2)$ (11)
-$Q_{1},Q_{2}=\text{Split}(\text{Shuffle}(Query),2)$
-Representation subspaces learning: The $Q_{1}$, $Q_{2}$ are reshaped into dimension $(\frac{C_{1}}{4},H\times W)$ and the $K_{1}$, $k_{2}$ are reshaped into dimension $(\frac{C_{1}}{4},\frac{H\times W}{f^{2}})$.
-Then, the two representation subspaces $W_{1}$ and $W_{2}$ are computed by dot product of $K$ and $Q$ respectively:
+It can be noted that if the vectors are directly split and self-similarity is calculated in different subspaces, it might result in the inability to simultaneously include RGB and Depth features in different subspaces.
+Therefore, we introduce a Shuffle mechanism to ensure each $K1$, $K2$, $Q1$, and $Q2$ contain information from both modalities.
+As shown in Figure 4.
+Figure 5: Splitting and Mixing of Multimodal Information.
+After channel shuffle, CMA can ensure that both RGB and Depth information can be utilized simultaneously when calculating self-similarity.
+3.3.4 Representation Learning in Multiple Subspaces
+Finally, CMA calculates the cross-modal similarity and embeds the calculation result into the fused feature $Value$, similar to operations in [31, 32, 8].
+Specifically, the study computes the dot products $K_{1}$, $Q_{1}$ and $K_{2}$, $Q_{2}$ to derive two representations of subspaces $W_{1}$ and $W_{2}$:
 $W_{1}=\text{Softmax}(\frac{Q_{1}\cdot K_{1}^{T}}{\sqrt{C_{1}/4}})$ (4)
 $W_{2}=\text{Softmax}(\frac{Q_{2}\cdot K_{2}^{T}}{\sqrt{C_{1}/4}})$
-The $W_{1}$ and $W_{2}$ are embedded into $V_{1}$ and $V_{2}$ by dot product. The fused feature $Fused_{2}$ is the concatenation of $V_{1}$ and $V_{2}$:
-$Fused_{2}=\text{Cat}[W_{1}\cdot V_{1},W_{2}\cdot V_{2}]\$ (13)
-Finally, the $Fused_{2}$ is adapted into output channel $C_{2}$ and added with residual connection $Fused_{1}$.
-$Fused_{2}=\text{ConvBN}_{1\times 1}(Fused_{2})+Fuesed_{1}$ (14)
-TABLE I: Ablation experiment results for different multi-modal feature fusion method.
+Then, $W_{1}$ and $W_{2}$ embed information into $V_{1}$ and $V_{2}$ through dot products. The fused feature $Fused_{2}$ is the concatenation of $V_{1}$ and $V_{2}$ in the channel dimension:
+$Fused=\text{Cat}[W_{1}\cdot V_{1},W_{2}\cdot V_{2}]\$ (5)
+Finally, CMA converts $Fused_{2}$ into output channels $C_{2}$ using a linear layer and connects it with the residual connection $Fused$, resulting in CMA’s output feature.
+Table 1: Ablation experiment results for different multi-modal feature fusion method.
 Model Feature Fusion Method Metric
@@
 33.0 (+1.1M) 54.1 (+7.1%) 65.5
-IV EXPERIMENT RESULTS AND ANALYSIS
-IV-A Implementation Details
-To evaluate our Real-Time semantic segmentation network design, we conduct a series of experiments on two widely-used datasets: NYUv2[5] (795 training and 654 testing RGB-D images) and SUNRGBD[6] (5825 training and 5050 testing RGB-D images).
+4 EXPERIMENT RESULTS
+Table 2: Comparison Results on NYUv2. The inference speed is tested on RTX 3090 platform, (480 $\times$ 640) inputs. MS denotes Multi-Scale inference strategy.
+Method Year Backbone Params/M mIoU (%) Real-Time Speed/FPS Speed (%)
+CMX-B2 [42] 2022 Segformer-B2 67 54.1 $\times$ 24 36.9%
+Token-Fusion [33] 2022 Token-Fusion(S) - 54.2 $\times$ - -
+CMX-B2 (MS) [42] 2022 Segformer-B2 67 54.4 $\times$ - -
+Multi-MAE [1] 2022 Vit-B - 56.0 $\times$ - -
+CMX-B4 (MS) [42] 2022 Segformer-B4 140 56.3 $\times$ - -
+Omnivore [10] 2022 Swin-L - 56.8 $\times$ - -
+CMX-B5 (MS) [42] 2022 Segformer-B5 181 56.9 $\times$ - -
+SA-Gate [5] 2021 Res50 65 50.2 $\checkmark$ 35 53.8%
+ESANet [26] 2022 Res34-Nbt1D 34 51.6 $\checkmark$ 32 49.2%
+PSCNet-L [9] 2022 Res50 52 46.2 $\checkmark$ 30 46.2%
+PSCNet-T [9] 2022 Res50 40 45.4 $\checkmark$ 47 72.3%
+PGDENet [49] 2022 Res34 101 53.7 $\checkmark$ 32 49.2%
+FRNet [50] 2022 Res34 86 53.6 $\checkmark$ 39 60.0%
+DFormer-S [37] 2023 DFormer-S 18.7 53.6 $\checkmark$ 50 76.9%
+AsymFormer 2024 B0+T 33 54.1 $\checkmark$ 65 100.0%
+AsymFormer (FP16) 2024 B0+T 33 54.1 $\checkmark$ 79 121.5%
+AsymFormer (MS) 2024 B0+T 33 55.3 $\times$ - -
+4.1 Implementation Details
+To evaluate our Real-Time semantic segmentation network design, we conduct a series of experiments on two widely-used datasets NYUv2[28] (795 training and 654 testing RGB-D images) and SUNRGBD[30] (5825 training and 5050 testing RGB-D images).
 We conduct the model training and testing on different platforms.
@@
 For the evaluation and inference speed testing, we use Nvidia RTX 3090 GPU, Ubuntu 20.04, CUDA 12.0 and Pytorch 2.0.1.
-We apply data augmentation to all datasets by randomly flipping (p=0.5), srandom scales [1.0,2.0], random crop 480$\times$640 and random HSV.
-We adopt Mix Transformer[17] and ConvNext[16] backbones pretrained on ImageNet-1k[24].
-In the augmented setting, both backbones are pretrained on ADE20k[23].
-For the CMA implementation, we set down-sample rate $f$=1 in CMA-1 and $f$=2 in CMA-2,3.
+We apply data augmentation to all datasets by randomly flipping (p=0.5), random scales between 1.0 and 2.0, random crop 480$\times$640 and random HSV.
+We adopt ConvNext-T [21] backbones pretrained on ImageNet-1k[6].
+For the Mix Transformer-B0 [35], no pre-trained weight has been used.
 The MLP-decoder in AsymFormer has the same structure as Segformer and an embedding dimension of 256.
 We choose AdamW optimizer with a weight decay of 0.01.
-The initial learning rate is $5e^{-5}$ and we use a poly learning rate schedule $(1-\frac{iter}{max_{i}ter})^{0.9r}$ with a warm-up of 10 epochs.
+The initial learning rate is $5e^{-5}$ and we use a poly learning rate schedule $(1-\frac{iter}{max_{i}ter})^{0.9}$ with a warm-up of 10 epochs.
 We train with a batch size of 8 for NYUv2 (500 epochs) and SUNRGBD (200 epochs).
@@
 The evaluation metric is mean Intersection over Union (mIoU).
-IV-B Model Assessment
-This paper proposes two different models based on the AsymFormer framework with different backbones.
-The AsymFormer uses ConvNext-T[16] for RGB representation learning and MiT-B0[17] for fused feature processing.
-This model has 33.0 million parameters and 36.0 GFLOPs computational cost, and it can achieve 65 FPS inference speed on RTX 3090.
-Particularly, with mixed precision quantization, AsymFormer achieves a real-time inference speed of 79 FPS on the RTX 3090 GPU.
-It also achieves a real-time inference speed of 29.4 FPS on the Tesla T4 GPU, which has a more limited computation resources (65 TFLOPs FP16).
-IV-C Ablation Experiment
+4.2 Ablation Experiment
 We conduct a series of ablation experiments on NYUv2 dataset to evaluate the effectiveness of the LAFS and CMA module.
@@
 Here, SE is used for feature selection in the channel direction, while MHSA is employed for further feature extraction on the fused features.
-Specifically, we use an MHSA with 2 heads to align with CMA.
 In our experiments, the Cat fusion method, used as a baseline, achieved a segmentation accuracy of 47.0 mIoU and an inference speed of 77.5 FPS.
-When using LAFS for feature selection alone, we achieved a performance improvement of 2.1% while sacrificing only 1.8 FPS of inference speed.
-This demonstrates that LAFS is a highly efficient feature selection method that provides performance gains without significantly impacting inference speed.
-In comparison to the other baseline, Cat+MHSA, which resulted in an 11.8 FPS reduction in inference speed, only a 2.9% improvement in segmentation accuracy was achieved.
+When using LAFS alone, we achieved a performance improvement of 2.1% while sacrificing only 1.8 FPS of inference speed.
+This demonstrates that LAFS provides performance gains without significantly impacting inference speed.
+In comparison to the other baseline, using Cat+MHSA, which resulted in a reduction of inference speed by 11.8 FPS, an improvement of only 2.9% in mIoU was achieved.
 This further highlights the efficiency of LAFS.
-Furthermore, we conducted experiments to investigate whether incorporating the mining of self-similarity within the two modalities after feature selection could further improve segmentation accuracy.
-When using CMA alone, we observed a 2.6% improvement in segmentation accuracy but encountered a significant decrease in inference speed of 10.1 FPS.
+Furthermore, when using CMA alone, we observed a 2.6% improvement in segmentation accuracy but encountered a significant decrease in inference speed of 10.1 FPS.
 Compared to LAFS, CMA showed a more noticeable reduction in inference speed.
-In comparison to SE+MHSA, CMA achieved similar accuracy, but it is important to note that no feature selection was performed in this case, indicating that relying solely on mining multi-modal self-similarity can achieve accuracy similar to the traditional feature fusion baseline SE+MHSA.
 Finally, we combined LAFS with CMA (LAFS+CMA).
-Since LAFS had minimal impact on inference speed and served a different purpose than CMA, the network’s inference speed only decreased by 2 FPS while achieving a significant improvement of 7.1% compared to the baseline Cat.
-At this point, the inference speed of LAFS+CMA was similar to SE+MHSA, but with a 2.1% performance improvement.
+Since LAFS had minimal impact on inference speed and served a different purpose than CMA, the network’s inference speed decreased by only 2 FPS.
+This change achieved a significant improvement of 7.1% in segmentation accuracy compared to the baseline Cat.
+At this point, the inference speed of LAFS+CMA was similar to SE+MHSA, but with a 4.2% performance improvement.
 This validates our experimental hypothesis: by re-modeling feature selection and mining cross-modal self-similarity, we can enhance the segmentation performance of the network without sacrificing inference speed compared to existing models.
 This demonstrates that we have indeed improved the efficiency of the network.
-IV-D Comparison With State-of-The-Arts
-IV-D1 NYUv2 Comparison Results
+4.3 Comparison With State-of-The-Arts
+4.3.1 NYUv2 Comparison Results
 […]
-TABLE II: Comparison Results on NYUv2. * denotes ADE20k pretrain. The inference speed is tested on RTX 3090 platform, (480 $\times$ 640) inputs.
-Method Year Backbone Params/M Pixel Acc. (%) mIoU (%) Inf Speed/FPS Inf Speed (%)
-PSCNet-T 2022 Res50 40 - 45.4 47 72.3%
-PSCNet-L 2022 Res50 52 - 46.2 30 46.2%
-ACNet 2019 Res50 118 - 48.3 28 43.1%
-RDFNet 2017 Res152 450 76 49.1 15 23.1%
-SA-Gate 2021 Res50 65 76.8 50.2 35 53.8%
-ESANet 2022 Res34-Nbt1D 34 - 51.6 32 49.2%
-MSFNet 2022 Res50 49 - 52.2 25 38.5%
-FRNet 2022 Res34 86 77.6 53.6 39 60.0%
-PGDENet 2022 Res34 101 78.1 53.7 32 49.2%
-CMX-B2 2022 Segformer-B2 67 78.7 54.1 24 36.9%
-AsymFormer 2023 B0+T 33 78.5 54.1 65 100.0%
-AsymFormer (Multi-Scale) 2023 B0+T 33 79.3 55.3 - -
-AsymFormer (FP16) 2023 B0+T 33 78.5 54.1 79 121.5%
-IV-D2 SUNRGBD Comparison Results
-Table III reports the performance of AsymFormer on the SUNRGBD dataset.
+4.3.2 SUNRGBD Comparison Results
+Table 3 reports the performance of AsymFormer on the SUNRGBD dataset.
 AsymFormer achieves competitive accuracy with 49.1 % mIoU.
-Since the input image size is the same as NYUv2, all methods have the same inference speed as shown in Table II.
-However, the advantage of AsymFormer is not as significant as in NYUv2 experiment.
+The advantage of AsymFormer is not as significant as in NYUv2 experiment.
 For example, AsymFormer improves 3.9 mIoU over SA-Gate[5] in NYUv2 dataset (54.1% vs 50.2% mIoU), but decreases 0.3 mIoU in SUNRGBD dataset (49.1% vs 49.4% mIoU).
@@
 Given that AsymFormer still has faster inference speed than other methods, we consider this performance acceptable for AsymFormer.
-TABLE III: Comparison Results on SUNRGBD. # denotes multi-scale test.
+Table 3: Comparison Results on SUNRGBD. MS denotes Multi-Scale inference strategy.
 Method Pixel Acc. (%) mIoU (%)
@@
 SA-Gate [5] 82.5 49.4
-CMX-B2[8] 82.8 49.7
+CMX-B2 (MS) [42] 82.8 49.7
+DFormer-S [37] - 50.0
 MSFNet [18] - 50.3
@@
 PGDENet [49] 87.7 51.0
+CMX-B4 (MS) [42] 83.5 52.1
+CMX-B5 (MS) [42] 83.8 52.4
 AsymFormer 81.9 49.1
-IV-E Visualization
-1) LAFS Attention Map: As shown in Figure 2, to demonstrate that LAFS performs better than CBAM [18] in selecting features in the spatial dimension, we visualized the spatial attention weights of both methods.
+4.4 Visualization
+4.4.1 LAFS Attention Map
+As shown in Figure 6, to demonstrate that LAFS performs better than CBAM [34] in selecting features in the spatial dimension, we visualized the spatial attention weights of both methods.
 It can be observed that LAFS provides better coverage of informative regions in the image while maintaining consistency within objects and preserving the integrity of edges.
-Figure 2: Difference between CBAM and LAFS’s spatial attention map.
-2) Semantic Segmentation Results: The Figure 3 demonstrates the segmentation results of AsymFormer on the NYUv2 dataset.
+Figure 6: Difference between CBAM and LAFS.
+4.4.2 Semantic Segmentation Results
+The Figure 7 demonstrates the segmentation results of AsymFormer on the NYUv2 dataset.
 As observed, while maintaining a significantly faster inference speed compared to other methods, AsymFormer achieves comparable semantic segmentation accuracy to mainstream approaches.
 Figure 7: Visualization of AsymFormer Semantic Segmentation Results.
-V CONCLUSIONS
-In this work, we proposed AsymFormer, which aims to construct a less-redundant real-time indoor scene understanding system.
+5 CONCLUSIONS
+In this work, we proposed AsymFormer, which aims to construct a less-redundant real-time indoor scene understanding network.
 To enhance efficiency and reduce redundant parameters, we implemented the following improvement: 1.
-We employed an asymmetric backbone that compressed the parameters of the Depth feature extraction branch, thus reducing redundancy.
+the asymmetric backbone that compressed the parameters of the Depth feature extraction branch, thus reducing redundancy.
 2.
-We introduced the LAFS module for feature selection, utilizing learnable feature weights to calculate spatial attention weights.
-This resulted in a significant performance improvement while almost maintaining the inference speed.
-3.
-Expanding on feature selection, we integrated modeling of self-similarity in multi-modal features, validating its capability to enhance network accuracy with minimal additional model parameters.
-The experiments demonstrated that the AsymFormer surpassed existing SOTA methods in terms of speed, demonstrating our method achieved a balance between accuracy and speed.
+the LAFS module for feature selection, utilizing learnable strategy for global information compressing and improving spatial attention calculations 3.
+the self-similarity in multi-modal features, validating its capability to enhance network accuracy with minimal additional model parameters.
+The experiments demonstrated that the AsymFormer achieves a balance between accuracy and speed.
 Moving forward, we will continue to optimize the modules and address issues such as self-supervised pre-training of the model, aiming for further improvements.